OpenStudy (anonymous):

Find the sum. Write in simplest form: 8 1/2+4 1/2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

13

OpenStudy (skullpatrol):

$$\Huge 8 \frac{1}{2}+4\frac{1}{2}$$$$\Huge =8 +\frac{1}{2}+4+\frac{1}{2}$$$$\Huge =8 +4+\frac{1}{2}+\frac{1}{2}$$$$\Huge =12+1$$ $$\Huge=13$$

OpenStudy (skullpatrol):

any questions @BabyyyGirll ?

OpenStudy (anonymous):

Nope Thanks @skullpatrol

OpenStudy (skullpatrol):

thanks for asking :)

OpenStudy (anonymous):

13